OpenStudy (anonymous):

integrate (6 sin x) (cos x + 1)^5 dx

OpenStudy (raden):

integral by u-sub let u=cos(x) + 1 du = -sin(x) dx or - du = sin(x)dx continue it ....

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

yeaa is the answer -(cos x + 1)^6 + C ?

OpenStudy (raden):

YOU ARE RIGHT :)

OpenStudy (anonymous):

thanks!

OpenStudy (raden):

you're welcome